OpenStudy (anonymous):

Solve the equation by using the LCD. Check for extraneous solutions. (4/x)+x=5 Please show work! Thanks.

OpenStudy (anonymous):

(4/x)+(x/1)=5 (4+x)/x=5 cross multiply 4+x=5x 4=5x-x 4=4x x=1

OpenStudy (anonymous):

I can't cross multiply, i'm have to use the LCD.... :( Thanks anyway but do you know how to use the LCD?

OpenStudy (anonymous):

i can type it here,but i can give u a link to show u how it's been solve on yahoo answers

OpenStudy (anonymous):

okay(: